# Convergent Primal-Dual Plug-and-Play Image Restoration: A General Algorithm and Applications

### v2

@@ I-B Plug-and-Play (PnP) Algorithms
PnP-FBS Pesquet et.al. [17] Quadric -- $J$ is firmly nonexpansive.
Ebner et al. [18] Quadric -- $J$ is contractive.
PnP-ADMM Venkatakrishnan et al. [19] $**{}^{\text{*}}$- -- -- -- --
Rond et al. [20] GKL div. -- -- -- -- --
Sreehari et al. [21] $**{}^{\text{*}}$- -- -- $\nabla J$ is a symmetric matrix, etc.
S. H. Chan et al. [22] $**{}^{\text{*}}$- -- -- $J$ is bounded, etc.
P. Nair et al. [23] $**{}^{\text{*}}$- -- -- $J$ is firmly nonexpansive.
Sun et al. [24] $**{}^{\text{*}}$- $J$ is firmly nonexpansive.
PnP-PDS Ono[25] $**{}^{\text{*}}$- -- --
@@
Following these impressive results, several works explore attempts to plug in more sophisticated denoisers. Zhang et al. proposed a neural network specifically designed for use in PnP methods [37], known as DRUNet. Duff et al. proposed generative regularizers [38], realized by Autoencoders (AE), Variational Autoencoders (VAE)[39], and Generative Adversarial Networks (GAN) [40]. Zhu et al. showed that diffusion models [41] also have promising performance as prior knowledge used in PnP methods [42]. We should mention that despite such extensive research, the theoretical convergence guarantee for PnP methods remains a challenging open question [43, 21, 22, 18]. This is mainly because replacing the proximity operator of the regularization term with a denoiser does not necessarily guarantee the existence of the underlying convex optimization problem.
To explore explicit data-driven regularization, alternative frameworks including regularization by denoising (RED) [44, 45], gradient step denoisers (GSD) [46, 47, 48, 49], and its generalization in Bregman geometry [50] have been proposed. Unlike PnP strategies, they explicitly define a regularization term derived from the output of a denoiser, incorporating it directly into the formulation of the optimization problem. Subsequently, they construct algorithms to solve this problem as a special case of traditional proximal splitting algorithms. Nonetheless, it should be noted that they still rely on conditions that may not be desirable in real-world settings. For example, RED imposes strict assumptions on the denoiser to guarantee its convergence, which are often impractical for deep denoisers [51] (see Section III-C4 for details). The GSD approach introduces a non-convex regularization term, which prevents guarantees of convergence to globally optimal solutions and makes its performance unreliable in practice.
In contrast, within the context of PnP methods, some studies have been conducted to establish their convergence under realistic assumptions, focusing on the firmly nonexpansiveness of denoisers. Pesquet et al. showed theoretically that a subclass of maximally monotone operators (MMO) can be fully represented by the firmly nonexpansive DnCNN with a certain network structure. They also constructed a PnP method based on forward-backward splitting (PnP-FBS) with convergence guarantees for image restoration [17]. As highlighted by the authors, their technique can be extended to any algorithm, as long as its analysis relies on the theory of MMOs, including PDS. Sun et al. proposed a PnP based on ADMM (PnP-ADMM) with the convergence guarantee by plugging in a firmly nonexpansive denoiser [24]. Hertrich et al. [52] developed a method to train averaged CNNs by stochastic gradient descent on the Stiefel manifold.
@@ II-A Notations
[…]
For $D\subset\mathcal{H}$, we consider an operator $T:D\to\mathcal{H}$. If $\|T\mathbf{x}_{1}-T\mathbf{x}_{2}\|\leq\|\mathbf{x}_{1}-\mathbf{x}_{2}\|$ holds for every $(\mathbf{x}_{1},\,\mathbf{x}_{2})\in D^{2}$, $T$ is said to be nonexpansive. When $2T-\mathrm{Id}$ is nonexpansive, $T$ is called firmly nonexpansive [54, Proposition 4.4] ($\mathrm{Id}$ denotes an identity operator). If $\kappa T$ is firmly nonexpansive for some $\kappa\in(0,\,\infty)$, then $T$ is $\kappa$-cocoercive.
@@
$\mathrm{GKL}_{\mathbf{v}}(\mathbf{x})$ (3)
$:=\displaystyle\sum_{i=1}^{N}\begin{cases}\eta[\mathbf{x}]_{i}-[\mathbf{v}]_{i}\ln{\eta[\mathbf{x}]_{i}},\,&\text{if $[\mathbf{v}]_{i}>0$ and $[\mathbf{x}]_{i}>0$},\\ \eta[\mathbf{x}]_{i},\,&\text{if $[\mathbf{v}]_{i}=0$ and $[\mathbf{x}]_{i}\geq 0$,}\\ \infty,\,&\text{otherwise,}\end{cases}$ (4)
@@ II-B Proximal Tools
According to the Moreau’s identity[56, Theorem 3.1 (ii)], $\mathrm{prox}_{\gamma\mathcal{F}^{*}}$ can be computed via $\mathrm{prox}_{\mathcal{F}/\gamma}$ as
@@ II-C Primal-Dual Splitting Algorithm
where the set of solutions to (11) is assumed to be nonempty. Its dual formulation is given as follows [54, Chap. 15]:
@@
$\begin{split}\mathrm{Find}\,&(\hat{\mathbf{x}},\,\hat{\mathbf{y}})\in\,\mathcal{X}\times\mathcal{Y}\\ \mathrm{s.t.}&\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\end{pmatrix}\in\begin{pmatrix}\partial g(\hat{\mathbf{x}})+\mathbf{L}^{*}\hat{\mathbf{y}}+\nabla f(\hat{\mathbf{x}})\\ -\mathbf{L}\hat{\mathbf{x}}+\partial h^{*}(\hat{\mathbf{y}})\end{pmatrix},\end{split}$ (15)
where ($\partial$($\cdot$)∂ ( ⋅ ) denotes the subdifferential of a convex function. Under a certain mild condition, the solutions to (15) coincide with the solutions to (11) and (12) (see [13] and [14] for details).
@@ II-D Firmly Nonexpansive Denoiser[17]
We review the method proposed in [17] for training a firmly nonexpansive denoiser represented by DnCNN, which is essential for the construction of convergent PnP algorithms. Basically, they add a specific penalty during training to ensure the firmly nonexpansiveness of the denoiser.
@@
By the definition of firmly nonexpansiveness, $J$ is firmly nonexpansive if and only if $Q$ is nonexpansive. Thus, our objective is to ensure the nonexpansiveness of $Q$ rather than the firmly nonexpansiveness of $J$. If $Q$ is differentiable, its nonexpansiveness is equivalent to its Jacobian $\nabla Q$ satisfying the following condition for all $\mathbf{x}\in\mathbb{R}^{K}$:
@@
[…]
@@
where $\rho$ is a random value chosen according to a uniform distribution in the range from 00 to $1$.
Note that $\|\nabla Q(\mathbf{\tilde{x}}_{\ell})\|_{\mathrm{sp}}$ can be computed using the power iterative method[57]. For more detailed information on the training method, see [17, Section 3.2].
@@ III Proposed Method
We propose a general convergent PnP-PDS, demonstrate its application to image restoration with specific cases of our PnP-PDS, and provide detailed discussions on its differences from related methods. For clarity, we present a flowchart in Fig 1 that illustrates the relationships between our theoretical results.
@@ III-A General Form of Convergent PnP-PDS and Its Convergence Analysis
$\begin{split}&\widetilde{\mathbf{x}}_{n+1}=J\left(\mathbf{x}_{n}-\gamma_{1}(\nabla f(\mathbf{x}_{n})+\mathbf{L}^{*}\mathbf{y}_{n})\right),\\ &\widetilde{\mathbf{y}}_{n+1}=\mathrm{prox}_{\gamma_{2}h^{*}}\left(\mathbf{y}_{n}+\gamma_{2}\mathbf{L}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\\ &\mathbf{x}_{n+1}=\rho_{n}\widetilde{\mathbf{x}}_{n+1}+(1-\rho_{n})\mathbf{x}_{n},\\ &\mathbf{y}_{n+1}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}+(1-\rho_{n})\mathbf{y}_{n},\end{split}$ (19)
where $\gamma_{1},\,\gamma_{2}$ and $\{\rho_{n}\}$ are the parameters and the definitions of $f$, $h$, and $\mathbf{L}$ are the same as in (11).
@@
$\begin{split}\mathrm{Find}\,&(\hat{\mathbf{x}},\,\hat{\mathbf{y}})\in\,\mathcal{X}\times\mathcal{Y}\\ &\mathrm{s.t.}\begin{pmatrix}\mathbf{0}\\ \mathbf{0}\end{pmatrix}\in\begin{pmatrix}\gamma_{1}^{-1}A_{J}(\hat{\mathbf{x}})+\mathbf{L}^{*}\hat{\mathbf{y}}+\nabla f(\hat{\mathbf{x}})\\ -\mathbf{L}\hat{\mathbf{x}}+\partial h^{*}(\hat{\mathbf{y}})\end{pmatrix}\end{split}$ (20)
@@ 3. $(iii)$
Recall that the set-valued operator $A_{J}=J^{-1}-\mathrm{Id}$ is maximally monotone if and only if $J$ is firmly nonexpansive and has the full domain [54, Proposition 23.8 (iii)]. Hence, $\gamma_{1}^{-1}A_{J}$ is also maximally monotone [54, Proposition 20.22]. Therefore, Proposition 3.1 and Proposition 3.2 are justified by [13, Theorem 3.1] and [13, Theorem 3.2], respectively. These proofs build upon and directly apply the standard convergence analysis of the PDS algorithm. For completeness and ease of reference, we provide the detailed proofs of these propositions in the Appendix. ∎
@@
The monotone inclusion problem in (20) has the same form as (15) with $\partial g=\gamma_{1}^{-1}A_{J}$. The classical monotone operator theory states that there exists $g\in\Gamma_{0}(\mathbb{R}^{N})$ such that $\partial g=\gamma_{1}^{-1}A_{J}$ if and only if $\gamma_{1}^{-1}A_{J}$ is a maximal cyclically monotone operator, a specific type of maximally monotone operator [54, Theorem 22.18]. However, the operator $\gamma_{1}^{-1}A_{J}$, where $J=(A_{J}+\mathrm{Id})^{-1}$, is not necessarily maximal cyclically monotone only because $J$ is firmly nonexpansive. Thus, the existence of $g\in\Gamma_{0}(\mathbb{R}^{N})$ cannot always be guaranteed. Furthermore, it should be noted that $A_{J}$ is scaled by $\gamma_{1}^{-1}$ in the solution set of the monotone inclusion problem (20). This implies that the value of $\gamma_{1}$ affects the solution set, unlike the standard PDS algorithm.
@@ III-B Applications to Image Restoration
$\begin{split}\underset{\mathbf{\hat{u}}\in\mathbb{R}^{K}}{\operatorname{minimize}}\,\mathcal{D}(\mathbf{\Phi}\mathbf{\hat{u}})+\mathcal{R}_{J}(\mathbf{\hat{u}})+\iota_{C}(\mathbf{\hat{u}}),\end{split}$ (24)
@@
$\begin{split}&\widetilde{\mathbf{x}}_{n+1}=J\left(\mathbf{x}_{n}-\gamma_{1}\left(\mathbf{\Phi}^{*}\mathbf{y}_{n}^{(1)}+\mathbf{y}_{n}^{(2)}\right)\right),\\ &\widetilde{\mathbf{y}}_{n+1}^{(1)}=\mathrm{prox}_{\gamma_{2}\mathcal{D}^{*}}\left(\mathbf{y}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\\ &\widetilde{\mathbf{y}}_{n+1}^{(2)}=\mathrm{prox}_{\gamma_{2}\iota_{C}^{*}}\left(\mathbf{y}_{n}^{(2)}+\gamma_{2}(2\mathbf{x}_{n+1}-\mathbf{x}_{n})\right),\\ &\mathbf{x}_{n+1}=\rho_{n}\widetilde{\mathbf{x}}_{n+1}+(1-\rho_{n})\mathbf{x}_{n},\\ &\mathbf{y}_{n+1}^{(1)}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}^{(1)}+(1-\rho_{n})\mathbf{y}_{n}^{(1)},\\ &\mathbf{y}_{n+1}^{(2)}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}^{(2)}+(1-\rho_{n})\mathbf{y}_{n}^{(2)},\end{split}$ (27)
@@
where $\mathbf{I}$ represents an identity matrix. Then, the monotone inclusion problem in (20) is reduced to (25) [54, Proposition 13.30, Proposition 16.9, Corollary 16.48], and we obtain Algorithm (27) from (19). The condition regarding $\gamma_{1}$ and $\gamma_{2}$ is obtained from the assumption in $(i)$ of Proposition 3.2 and the following fact [58, Theorem 1]:

### v5

@@ I-B Plug-and-Play (PnP) Algorithms
PnP-FBS Sun et.al. [17] Quadratic - - $J$ is averaged.
Pesquet et.al. [18] Quadratic - $J$ is firmly nonexpansive.
Ebner et al. [19] Quadratic - $J$ is contractive.
[…]
Rond et al. [21] GKL div. - - - - -
[…]
@@
Following these impressive results, several works explore attempts to plug in more sophisticated denoisers. Zhang et al. proposed a neural network specifically designed for use in PnP methods [38], known as DRUNet. Duff et al. proposed generative regularizers [39], realized by Autoencoders (AE), Variational Autoencoders (VAE) [40], and Generative Adversarial Networks (GAN) [41]. Zhu et al. showed that diffusion models [42] also have promising performance as prior knowledge used in PnP methods [43]. We should mention that despite such extensive research, the theoretical convergence guarantee for PnP methods remains a challenging open question [44, 22, 23, 17, 19]. This is mainly because replacing the proximity operator of the regularization term with a denoiser does not necessarily guarantee the existence of the underlying convex optimization problem.
To explore explicit data-driven regularization, alternative frameworks including regularization by denoising (RED) [45, 46], gradient step denoisers (GSD) [47, 48, 49, 50], and its generalization in Bregman geometry [51] have been proposed. Unlike PnP strategies, they explicitly define a regularization term derived from the output of a denoiser, incorporating it directly into the formulation of the optimization problem. Subsequently, they construct algorithms to solve this problem as a special case of traditional proximal splitting algorithms. Nonetheless, it should be noted that they still rely on conditions that may not be desirable in real-world settings. For example, RED imposes strict assumptions on the denoiser to guarantee its convergence, which are often impractical for deep denoisers [52] (see Section III-C4 for details). The GSD approach introduces a non-convex regularization term, which makes it difficult to guarantee convergence to globally optimal solutions and may cause the final solution to depend on the initialization.
In contrast, within the context of PnP methods, some studies have been conducted to establish their convergence under realistic assumptions, focusing on the firm nonexpansiveness of denoisers. Pesquet et al. showed theoretically that a subclass of maximally monotone operators (MMO) can be fully represented by the firmly nonexpansive DnCNN with a certain network structure. They also constructed a PnP method based on forward-backward splitting (PnP-FBS) with convergence guarantees for image restoration [18]. As highlighted by the authors, their technique can be extended to any algorithm, as long as its analysis relies on the theory of MMOs, including PDS. Sun et al. proposed a PnP based on ADMM (PnP-ADMM) with the convergence guarantee by plugging in a firmly nonexpansive denoiser [25]. Hertrich et al. [53] developed a method to train averaged CNNs by stochastic gradient descent on the Stiefel manifold.
@@ II-A Notations
Let $\mathcal{H}$ be a real Hilbert space with its inner product $\langle\cdot,\cdot\rangle$ and norm $\|\cdot\|$. The notation $\Gamma_{0}(\mathcal{H})$ represents the set of proper lower-semicontinuous convex functions from $\mathcal{H}$ to $\mathbb{R}\cup\{\infty\}$. For $\mathcal{F}\in\Gamma_{0}(\mathcal{H})$, its effective domain is $\mathrm{dom}(\mathcal{F}):=\{\mathbf{x}\in\mathcal{H}:\mathcal{F}(x)<\infty\}$.
For $D\subset\mathcal{H}$, we consider an operator $T:D\to\mathcal{H}$. If $\|T\mathbf{x}_{1}-T\mathbf{x}_{2}\|\leq\|\mathbf{x}_{1}-\mathbf{x}_{2}\|$ holds for every $(\mathbf{x}_{1},\,\mathbf{x}_{2})\in D^{2}$, $T$ is said to be nonexpansive. When $2T-\mathrm{Id}$ is nonexpansive, $T$ is called firmly nonexpansive [55, Proposition 4.4] ($\mathrm{Id}$ denotes an identity operator). If $\kappa T$ is firmly nonexpansive for some $\kappa\in(0,\,\infty)$, then $T$ is $\kappa$-cocoercive.
@@
$\hskip-11.38109pt\mathrm{GKL}_{\mathbf{v}}(\mathbf{x})$ (3)
$\hskip-11.38109pt:=\displaystyle\sum_{i=1}^{N}\begin{cases}\eta[\mathbf{x}]_{i}-[\mathbf{v}]_{i}\ln{\eta[\mathbf{x}]_{i}},\,&\hskip-5.69054pt\text{if $[\mathbf{v}]_{i}>0$ and $[\mathbf{x}]_{i}>0$},\\ \eta[\mathbf{x}]_{i},\,&\hskip-5.69054pt\text{if $[\mathbf{v}]_{i}=0$ and $[\mathbf{x}]_{i}\geq 0$,}\\ \infty,\,&\hskip-5.69054pt\text{otherwise,}\end{cases}$ (4)
@@ II-B Proximal Tools
According to the Moreau’s identity[57, Theorem 3.1 (ii)], $\mathrm{prox}_{\gamma\mathcal{F}^{*}}$ can be computed via $\mathrm{prox}_{\mathcal{F}/\gamma}$ as
@@ II-C Primal-Dual Splitting Algorithm
where the set of solutions to (11) is assumed to be nonempty. Its dual formulation is given as follows [55, Chap. 15]:
@@
[…]
where $\partial(\cdot)$ denotes the subdifferential of a convex function. Under a certain mild condition, the solutions to (15) coincide with the solutions to (11) and (12) (see [13] and [14] for details).
@@ II-D Firmly Nonexpansive Denoiser[18]
We review the method proposed in [18] for training a firmly nonexpansive denoiser represented by DnCNN, which is essential for the construction of convergent PnP algorithms. Basically, they add a specific penalty during training to ensure the firm nonexpansiveness of the denoiser.
@@
By the definition of firm nonexpansiveness, $J$ is firmly nonexpansive if and only if $Q$ is nonexpansive. Thus, our objective is to ensure the nonexpansiveness of $Q$ rather than the firm nonexpansiveness of $J$. If $Q$ is differentiable, its nonexpansiveness is equivalent to its Jacobian $\nabla Q$ satisfying the following condition for all $\mathbf{x}\in\mathbb{R}^{K}$:
@@
$\|J\left(\mathbf{x}_{\ell})-\overline{\mathbf{x}}_{\ell}\|^{2}_{2}+\tau\,\mathrm{max}(\|\nabla Q(\mathbf{\tilde{x}}_{\ell})\|_{\mathrm{sp}}^{2},\,1-\xi\right).$ (17)
@@
where $\rho$ is a random value chosen according to a uniform distribution in the range from 0 to $1$.
Note that $\|\nabla Q(\mathbf{\tilde{x}}_{\ell})\|_{\mathrm{sp}}$ can be computed using the power iterative method[58]. For more detailed information on the training method, see [18, Section 3.2].
@@ III Proposed Method
We propose a general convergent PnP-PDS, demonstrate its application to image restoration with specific cases of our PnP-PDS, and provide detailed discussions on its differences from related methods. For clarity, we present a flowchart in Fig. 1 that illustrates the relationships between our theoretical results.
@@ III-A General Form of Convergent PnP-PDS and Its Convergence Analysis
$\begin{split}&\widetilde{\mathbf{x}}_{n+1}=J\left(\mathbf{x}_{n}-\gamma_{1}(\nabla f(\mathbf{x}_{n})+\mathbf{L}^{*}\mathbf{y}_{n})\right),\\ &\widetilde{\mathbf{y}}_{n+1}=\mathrm{prox}_{\gamma_{2}h^{*}}\left(\mathbf{y}_{n}+\gamma_{2}\mathbf{L}(2\widetilde{\mathbf{x}}_{n+1}-\mathbf{x}_{n})\right),\\ &\mathbf{x}_{n+1}=\rho_{n}\widetilde{\mathbf{x}}_{n+1}+(1-\rho_{n})\mathbf{x}_{n},\\ &\mathbf{y}_{n+1}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}+(1-\rho_{n})\mathbf{y}_{n},\end{split}$ (19)
where $\gamma_{1},\,\gamma_{2}$ and $\{\rho_{n}\}$ are the parameters and the definitions of $f$, $h$, and $\mathbf{L}$ are the same as in (11). The parameters $\gamma_{1}$ and $\gamma_{2}$ are fixed constants across the iterations.
@@
[…]
@@
Recall that the set-valued operator $A_{J}=J^{-1}-\mathrm{Id}$ is maximally monotone if and only if $J$ is firmly nonexpansive and has the full domain [55, Proposition 23.8 (iii)]. Hence, $\gamma_{1}^{-1}A_{J}$ is also maximally monotone [55, Proposition 20.22]. Therefore, Proposition 3.1 and Proposition 3.2 are justified by [13, Theorem 3.1] and [13, Theorem 3.2], respectively. These proofs build upon and directly apply the standard convergence analysis of the PDS algorithm. For completeness and ease of reference, we provide the detailed proofs of these propositions in the Appendix. ∎
@@
The monotone inclusion problem in (20) has the same form as (15) with $\partial g=\gamma_{1}^{-1}A_{J}$. The classical monotone operator theory states that there exists $g\in\Gamma_{0}(\mathbb{R}^{N})$ such that $\partial g=\gamma_{1}^{-1}A_{J}$ if and only if $\gamma_{1}^{-1}A_{J}$ is a maximal cyclically monotone operator, a specific type of maximally monotone operator [55, Theorem 22.18]. However, the operator $\gamma_{1}^{-1}A_{J}$, where $J=(A_{J}+\mathrm{Id})^{-1}$, is not necessarily maximal cyclically monotone only because $J$ is firmly nonexpansive. Thus, the existence of $g\in\Gamma_{0}(\mathbb{R}^{N})$ cannot always be guaranteed. Furthermore, it should be noted that $A_{J}$ is scaled by $\gamma_{1}^{-1}$ in the solution set of the monotone inclusion problem (20). This implies that the value of $\gamma_{1}$ affects the solution set, unlike the standard PDS algorithm.
@@ III-B Applications to Image Restoration
[…]
@@
$\begin{split}&\widetilde{\mathbf{x}}_{n+1}=J\left(\mathbf{x}_{n}-\gamma_{1}\left(\mathbf{\Phi}^{*}\mathbf{y}_{n}^{(1)}+\mathbf{y}_{n}^{(2)}\right)\right),\\ &\widetilde{\mathbf{y}}_{n+1}^{(1)}=\mathrm{prox}_{\gamma_{2}\mathcal{D}^{*}}\left(\mathbf{y}_{n}^{(1)}+\gamma_{2}\mathbf{\Phi}(2\widetilde{\mathbf{x}}_{n+1}-\mathbf{x}_{n})\right),\\ &\widetilde{\mathbf{y}}_{n+1}^{(2)}=\mathrm{prox}_{\gamma_{2}\iota_{C}^{*}}\left(\mathbf{y}_{n}^{(2)}+\gamma_{2}(2\widetilde{\mathbf{x}}_{n+1}-\mathbf{x}_{n})\right),\\ &\mathbf{x}_{n+1}=\rho_{n}\widetilde{\mathbf{x}}_{n+1}+(1-\rho_{n})\mathbf{x}_{n},\\ &\mathbf{y}_{n+1}^{(1)}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}^{(1)}+(1-\rho_{n})\mathbf{y}_{n}^{(1)},\\ &\mathbf{y}_{n+1}^{(2)}=\rho_{n}\widetilde{\mathbf{y}}_{n+1}^{(2)}+(1-\rho_{n})\mathbf{y}_{n}^{(2)},\end{split}$ (27)
@@
where $\mathbf{I}$ represents an identity matrix. Then, the monotone inclusion problem in (20) is reduced to (25) [55, Proposition 13.30, Proposition 16.9, Corollary 16.48], and we obtain Algorithm (27) from (19). The condition regarding $\gamma_{1}$ and $\gamma_{2}$ is obtained from the assumption in (i) of Proposition 3.2 and the following fact [59, Theorem 1]: